OpenStudy (anonymous):

the point -3,2 and has a slope of three write this in point slope form slope intercept form and standard form plz help me

OpenStudy (anonymous):

When you have a point (x1,y1) and slope (m) point slope form is y-y1=m(x-x1) so plug your point and slope into that

OpenStudy (anonymous):

the slope would be m but where do i put the point at

OpenStudy (anonymous):

for x1 and y1 (-3,2) ^ ^ x1 y1

OpenStudy (anonymous):

so it would be y-2=3(x-3)

OpenStudy (anonymous):

Close, gotta change the sign because plugging in the x we have y-2=3(x-(-3) two negatives make a positive y-2=3(x+3)

OpenStudy (anonymous):

so thats your point slope form were done with that part

OpenStudy (anonymous):

ok thanks man im just trying to ace a test im confused on can you help me with the other two forms

OpenStudy (anonymous):

Yep, to get slope intercept form, just take point slope form and multiply out the right hand side and then isolate y

OpenStudy (anonymous):

what would you multiply by

OpenStudy (anonymous):

just multiply the parentheses y-2=3(x+3) y-2=3x+9 understand so far?

OpenStudy (anonymous):

yeah then whats next after that

OpenStudy (anonymous):

isolate y by adding 2 to both sides y=3x+11 and thats your slope intercept form

OpenStudy (anonymous):

then how do you get the last form

OpenStudy (anonymous):

Standard form looks like ax+by=c so we wanna get x and y on the left side in that order and the constant on the right subtract 3x from both sides -3x+y=11 looks good but sometimes they want the leading coefficient to be positive so we can make that happen my multiplying the entire equation by -1 3x-y=-11

OpenStudy (anonymous):

so 3x-y=-11 is the standard form so it was that easy

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

well thanks man i cant tell you how much of a help you were i really appreciate it

OpenStudy (anonymous):

You're welcome